OpenStudy (anonymous):

integral x^3/sqrt(1+x^4)

OpenStudy (campbell_st):

well its looks a bit like the numerator is the derivative of the denominator to me... what would that indicate...

OpenStudy (campbell_st):

as a hint whats the derivative of \[y = (1 = x^4)^{\frac{1}{2}}\]

OpenStudy (yttrium):

you can use substitution here let u = (1+x^4) du = 4x^3dx therefore x^3dx = du/(4) from there, you can see that \[\frac{ 1 }{ 4 } \int\limits \frac{ du }{ u^{\frac{ 1 }{ 2 }} }\] Then apply the integration principle. but take note that u = 1+x^4

OpenStudy (anonymous):

thank u Yttrium but where I'm having trouble is finding the integration principle can u help me?

OpenStudy (yttrium):

do you know power rule?

OpenStudy (anonymous):